Chocolate coffee beans sell for $7.00 per pound and hazelnut coffee beans sell for $6.10 per pound. One customer wants a 6 pound mixture of both types of coffee. How many pounds of each should be used if the mixture is to cost $6.40 per pound?
I have tried making tables,and writing an equation from each, but I keep getting the wrong answer and I am very confused. I need to know how to set these up to solve them, then I will be fine. It is the writing of the initial equation that I can't figure out.
Thanks for your help.

Answer by LilSkittleMd(119)

Let c=chocolate coffee
Let h=hazelnut coffee
Since you need a 6 pound mixture of both coffees, you have to add them together, so your first equation will be:
c+h=6
You need a 6 pound mixture that costs 6.40 per pound. For this equation you take the price per coffee multiplied by the type of coffee. And since you need a 6 pound mixture that costs 6.40 per pound, multipy 6.40 by 6 to get the total of the mixture.
7c+6.10h=38.4
From there you can solve by substitution.
c=6-h
7(6-h)+6.1h=38.4
42-7h+6.1h=38.4
42-0.9h=38.4
-0.9h=-3.6
h=4
c+4=6
c=2
2 pounds of chocolate coffee should be used
4 pounds of hazelnut coffee should be used.
Hope that helps.